# Mechanistic–statistical inference of mosquito dynamics from mark–release–recapture data

### v1

@@ Abstract
Biological control strategies against mosquito-borne diseases—such as the sterile insect technique (SIT), RIDL, and Wolbachia-based releases—require reliable field estimates of dispersal and survival of released males. We develop a mechanistic–statistical framework tailored to mark–release–recapture (MRR) data that links two complementary models: (i) a microscopic individual-based model (2D Itô diffusion with stochastic death and capture), and (ii) its macroscopic reaction–diffusion limit, which provides expected population densities and trap-specific capture intensities. Inference proceeds by solving the reaction–diffusion system and embedding it in a Poisson observation model for daily trap counts; uncertainty is quantified with a parametric bootstrap based on the individual-based model. We first validate parameter identifiability using simulated data that closely mimic the experimental data. We then analyze an urban MRR campaign in El Cano (Havana, Cuba) comprising four weekly releases of differentially marked sterile Aedes aegypti males and a network of indoor BG-Sentinel traps. A homogeneous mobility model is favored by the Akaike Information Criterion (AIC) over an urban–nonresidential heterogeneous alternative. Estimates indicate a post-release life expectancy of about five days ($\approx$ eight days total adult age) and a typical displacement on the order of 180 m after five days, with trap efficiency jointly identified alongside movement and survival. Unlike empirical analyses that summarize trap returns (e.g., exponential fits of daily survival or mean recapture distance), the mechanistic formulation couples movement, mortality, and capture explicitly and reduces biases from trap layout—enabling joint estimation and biologically interpretable parameters. The approach delivers a calibrated, computationally efficient procedure for extracting biologically interpretable parameters from sparse MRR data and offers a principled alternative to empirical summaries for the design and evaluation of SIT-style interventions in urban settings.
@@ 1 Introduction
Mosquitoes are the primary arthropod vectors of human diseases worldwide, transmitting numerous illnesses including malaria, lymphatic filariasis, and arboviruses such as dengue, chikungunya and Zika virus. For several of these diseases, available treatments are only symptomatic, and for most, no widely deployable specific vaccines or drugs currently exist. Under these circumstances, controlling mosquito populations remains the primary means of prevention. Given the limitations of conventional insecticide-based approaches, there is an increasing need for innovative strategies that are both sustainable and environmentally friendly ACH19 , BEC20 . Among these alternatives, biological control methods, such as the sterile insect technique (SIT), the release of insects carrying a dominant lethal (RIDL), or Wolbachia-based strategies, involve the release of large numbers of mosquitoes that are either sterile or less capable of transmitting disease.
Aedes aegypti is the primary vector of dengue worldwide. This mosquito lives in close association with humans and shows a strong preference for feeding on human blood. Its larvae can develop in a wide range of water reservoirs associated with household activities HAM07 . These characteristics make it an ideal vector for the spread of dengue virus, particularly in large cities with dense populations and numerous artificial water containers. In this context, the SIT has been progressively evaluated—from laboratory studies to large-cage experiments—by one of the present authors and collaborators, to systematically assess potential impacts on mosquito performance and survival under increasingly natural conditions GAT14 , GAT21 . However, a critical gap remains: reliable estimation of key ecological parameters of released sterile males (such as dispersal and survival) in real-life urban settings, which are essential to the design and scheduling of release strategies.
@@
MRR experiments, combined with spatio-temporal dispersal models, have long been used to infer demographic parameters including survival rates, longevity, and emigration MRR08 ; GroGim10 ; LebBur92 ; SIL07 . Yet, mosquito MRRs pose distinctive challenges: short lifespans and low recapture rates yield sparse, noisy data GAR16 , and classical population-scale models can be inadequate. This motivates the use of microscopic individual-based models (IBMs), which represent each mosquito explicitly, but IBMs are computationally intensive and their stochasticity complicates parameter inference HarCal11 .
Our methodological contribution is to integrate two complementary modeling frameworks. At the microscopic level, we model mosquito movement as a two-dimensional Itô diffusion with spatially varying mobility, and we treat death and capture as stochastic events. At the macroscopic level, we derive a reaction–diffusion formulation that is the deterministic limit of the IBM and yields the expected population density and expected captures. This theoretically grounded link allows us to estimate IBM parameters without direct large-scale simulation: inference proceeds by solving the reaction–diffusion system and embedding it in a mechanistic–statistical observation model. Such mechanistic–statistical approaches have proved effective for reaction–diffusion systems under sparse, noisy observations—either when models are constructed from biological principles RoqBon16 ; RoqSouRou11 or when they summarize underlying stochastic processes RoqSou22 ; RoqWal16 —and naturally fit within a state-space framework DurKoo12 ; PatTho08 . A central focus here is the explicit modeling of trapping and the joint identifiability of movement, survival, and trap efficiency.
@@
We remark that obtaining good estimates on how mosquito spatial diffusion depends on the heterogeneity of the terrain is important to be able to calibrate mathematical models that take into account spatially heterogeneous environments in order to control mosquito populations in an efficient and robust way (see, for instance, AgboAlmeidaCoron2 ).
@@ 2.1 Mark-release-recapture data
The mark-release-recapture (MRR) experiment described in this paper was conducted at a time when wild A. aegypti mosquito populations at the study site exhibited low abundance due to a preceding SIT pilot study GAT21 .
@@
Study Design: The study was designed following the guidelines described by Bouyer24 . Four releases of 10,000 marked mosquitoes each were carried out over consecutive weeks at a single location, referred to as the release point, at the center of the study site. Each week, the mosquitoes were marked with a different color using a fluorescent powder: yellow, red, blue, and pink, respectively. An adult mosquito trap network was deployed to recapture the released mosquitoes and monitor the wild population. The network consisted of twenty-one BG-Sentinel traps baited with BG-lures (Biogents, Germany), distributed along concentric rings at distances of approximately 50, 100, 150, 200, 250, 300 and 400 meters. The traps were installed indoors, at ground level, in the quietest areas of inhabited buildings, and were checked daily for six weeks. The collected biological material was transported to the laboratory (at IPK) in plastic containers, where insects were killed by freezing at -20°C. Mosquitoes were then identified morphologically by species and sex under a stereomicroscope. Males were further classified as either wild, unmarked individuals or marked with different dust colors, using ultraviolet light for identification.
Mosquitoes: A. aegypti were colonized from eggs collected in ovitraps at the study site in 2018. Adults were reared in cages at a density of one mosquito/cm², with a 1:2 male-to-female ratio, and fed a 10% honey solution. The colony was maintained at 28 ± 2°C and 80 ± 10% relative humidity. Porcine blood at 38°C was provided in collagen casings (Fibran, Girona, Spain) once a week for female feeding. Eggs were collected 2–3 days post female feeding and allowed to mature in a wet environment for 3 days. Larvae were reared in 100 cm × 60 cm × 3 cm trays, each containing 4 L of deionized water, at a density of 2 larvae/mL. Larval density was determined using egg quantity-weight regression curves as described in zheng2015standard . Simultaneous hatching was induced by immersing eggs in 36°C de-oxygenated water under vacuum. First-instar larvae were then transferred to trays. The IAEA standard diet (50% tuna meal, 36% bovine liver powder, and 14% brewer’s yeast) was provided daily at the rates of 0.2, 0.4, 0.8, and 0.6 mg/larva for larval instars I, II, III, and IV, respectively Puggioli2013 . After the onset of pupation, immatures were collected and sorted by sex using a Fay-Morlan apparatus (John W. Hock Co, Gainesville, FL, USA). Male pupae were dosed by volume into 10 mL plastic tubes, each containing approximately 500 individuals. Batches of 5,000 male pupae were kept in 1 L flat tissue-culture flasks (Thermo Fisher, Waltham, MA, USA) filled with 250 mL of dechlorinated water until the optimal age for irradiation was reached. The flasks were placed horizontally to maximize the water surface area.
@@
Marking: The boxes containing two-day-old adult sterile males were placed in a fridge at 4°C for 15 minutes to immobilize the mosquitoes. The immobilized mosquitoes were transferred in batches of approximately 2,500 specimens to 1-liter plastic containers containing 10.4 mg of fluorescent powder (DayGlo® Color Corp., USA) following previously described procedures Bouyer24 . The containers were gently rotated for ten seconds to ensure all mosquitoes came into contact with the pigment. Marked mosquitoes were then transferred to 30 x 30 x 30 cm metallic cages (BioQuip, USA). Mosquitoes were provided with water and honey solution for 24 hours prior to release.
@@ 2.2 Mechanistic models and simulated data
where $dB_{t}$ denotes an increment in time of a standard 2-dimensional Brownian motion Gar09 , and $\sigma$ is a Lipschitz-continuous function on $\mathbb{R}^{2}$ that describes the local mosquito mobility, which may vary depending on local conditions.
@@
First, ignoring the dead or alive status of the mosquitoes, the probability density $v(t,x)$ of the random variable $X_{t}$ is given by the standard Fokker-Planck equation Gar09 :
@@
Let $p(t,x)$ denote the probability density function of the mosquito being alive at time $t$ and located at $x$. The killing mechanism removes probability mass at a rate $F(x)p(t,x)$. Thus, the corresponding Fokker-Planck equation is given by (see, e.g., RoqSou22 ):
@@ 2.4 Numerical and software aspects
The 2D MRR simulator is implemented in Python and builds upon the MSE software , which is dedicated to the simulation and calibration of PDE models. The MRR simulator is available in a dedicated Git repository, here [https://forge.inrae.fr/msegrp/mrr-simulator]. Since MSE is distributed through GUIX courtes:hal-04586520 , this work can be reproduced in a controlled environment on Linux operating systems.
@@ 3.2 Parameter estimation using experimental data
To better compare these two models, we use the Akaike Information Criterion (AIC), which introduces a penalty term for the number of parameters in the model (Aka74, ): $AIC=-2\ln(\mathcal{L}(\Theta^{*}))+2N_{p},$ where $\mathcal{L}(\Theta^{*})$ is the maximum likelihood of the model, and $N_{p}$ is the number of parameters in the model ($N_{p}=3$ in the homogeneous model, $N_{p}=4$ in the heterogeneous model). The AIC balances the goodness of fit (measured by $2\ln(\mathcal{L}(\Theta^{*}))$) and model complexity (penalized by $2N_{p}$). A lower AIC value indicates a better model.
@@ 4 Discussion
Classical analyses of MRR datasets often summarize dispersal and survival empirically from trap returns (e.g., exponential fits for daily survival, recapture-weighted averages for dispersal, MuiKay98 ). In contrast, our approach links movement, mortality, and capture through explicit mechanisms and computes likelihoods from model-predicted captures at each trap and day, which avoids ad hoc transformations and, crucially, clarifies the interpretation of dispersal metrics. Our estimates are consistent with GAT21 for lifespan: in that other Cuban field trial, the probability of daily survival (PDS) was obtained by fitting an exponential model to recapture counts, and the average post-release life expectancy was computed as $1/(-\ln\mathrm{PDS})$, yielding 3.76 days; combined with the $\sim$3-day age at release, this corresponds to a total lifespan close to 7 days, in good agreement with our inference (about 8 days). For context, published field MRR studies of wild (non-sterile) A. aegypti males report daily survival probabilities around 0.570.570.700.70 in northern Australia MuiKay98 (implying mean life expectancy of $\approx 1.8$$2.8$ days) and $\approx 0.77$ on the Kenyan coast McD77I (implying $\approx 3.8$ days). Regarding pre-release age, MuiKay98 report that released males were 1–2 days post-emergence, whereas McD77I describes “newly emerged” males without specifying an exact number of days. These values place our sterile-male post-release estimate ($5$ days) and total lifespan (8 days) slightly above these wild-male estimates.
@@
By contrast, dispersal metrics differ substantially. In GAT21 , dispersal was summarized from the locations of recaptured marked males in a BG-Sentinel network arranged in concentric rings up to 400 m, as a mean distance of recaptures (77.3 m) and as flight ranges (43.2 m for 50% of recaptures, 110.5 m for 90%). Methodologically, this “mean distance” is the average of trap-to-release distances weighted by how many mosquitoes were recaptured in each trap; in other words, it is the mean distance conditional on being captured (this is a standard approach, also used in McD77II ; MuiKay98 ). Such a quantity is generally a downward-biased proxy for the population mean displacement because (i) capture probability typically declines with distance as trap density per unit circumference decreases with increasing ring radius, although this effect can be corrected for by applying a weighting factor, as proposed in MuiKay98 ; (ii) long-distance movements are right-truncated by the outer trap radius (400 m); and (iii) capture is an absorbing event that truncates trajectories—individuals caught early cannot contribute longer displacements they might otherwise have achieved. For these reasons, the reported distances should be interpreted as lower bounds on true dispersal. In our mechanistic–statistical framework, dispersal is inferred through an explicit diffusion parameter ($\sigma$), independent of trap placement. With the homogeneous-model estimate $\sigma=64~\text{m}/\sqrt{\text{day}}$, the expected distance from the release point is $E(\|X_{t}\|)=\sigma\sqrt{\pi t/2}\approx 80\sqrt{t}$ m, i.e., roughly 180 m after 5 days, naturally exceeding the uncorrected mean of recaptures and aligning with the expectation that trap-weighted averages under-estimate long-range movement.
@@
Regarding potential improvements to the MRR experiment, a promising alternative would be to perform simultaneous releases of mosquitoes marked with distinct colors at different locations, rather than successive weekly releases from the same site. In such a design, frequency data (i.e., the relative proportions of each color captured per trap) could be analyzed instead of absolute abundances, as proposed in RoqWal16 . Using proportions has the advantage of reducing sensitivity to heterogeneity in trapping efficiency across sites (e.g., spatial variations in $\gamma$ or $R$), thereby mitigating potential biases in parameter estimation.

### v2

@@ Abstract
Biological control strategies against mosquito-borne diseases require reliable field estimates of dispersal and survival of released males. We present a mechanistic–statistical framework for mark–release–recapture (MRR) data that links (i) an individual-based 2D Itô diffusion model with stochastic death and capture, and (ii) its reaction–diffusion approximation yielding expected densities and trap-specific captures. Inference is achieved by solving the reaction–diffusion system within a Poisson observation model for daily trap counts, with uncertainty assessed via parametric bootstrap. We first validate parameter identifiability using simulated data that closely mimic the experimental data. We then analyze an urban MRR campaign in El Cano (Havana, Cuba) comprising four weekly releases of differentially marked sterile Aedes aegypti males and a network of indoor BG-Sentinel traps. A homogeneous mobility model is favored by AIC over an urban–nonresidential heterogeneous alternative. Estimates indicate a post-release life expectancy of about five days (eight days total adult age) and a typical displacement of 180 m after five days, with trap efficiency jointly identified alongside movement and survival. Unlike empirical analyses that summarize trap returns (e.g., exponential fits of daily survival or mean recapture distance), the mechanistic formulation couples movement, mortality, and capture explicitly and reduces biases from trap layout—enabling joint estimation and biologically interpretable parameters. The approach delivers a calibrated, computationally efficient procedure for extracting biologically interpretable parameters from sparse MRR data and offers a principled alternative to empirical summaries for the design and evaluation of sterile insect technique (SIT)-style interventions in urban settings.
@@ 1 Introduction
Mosquitoes are the primary arthropod vectors of human diseases worldwide, transmitting numerous illnesses including malaria, lymphatic filariasis, and arboviruses such as dengue, chikungunya and Zika virus. For several of these diseases, available treatments are only symptomatic, and for most, no widely deployable specific vaccines or drugs currently exist. Under these circumstances, controlling mosquito populations remains the primary means of prevention. Given the limitations of conventional insecticide-based approaches, there is an increasing need for innovative strategies that are both sustainable and environmentally friendly (Achee et al., 2019; Becker et al., 2020). Among these alternatives, biological control methods, such as the sterile insect technique (SIT), the release of insects carrying a dominant lethal (RIDL), or Wolbachia-based strategies, involve the release of large numbers of mosquitoes that are either sterile or less capable of transmitting disease.
Aedes aegypti is the primary vector of dengue worldwide. This mosquito lives in close association with humans and shows a strong preference for feeding on human blood. Its larvae can develop in a wide range of water reservoirs associated with household activities (Hammond et al., 2007). These characteristics make it an ideal vector for the spread of dengue virus, particularly in large cities with dense populations and numerous artificial water containers. In this context, the SIT has been progressively evaluated—from laboratory studies to large-cage experiments—by one of the present authors and collaborators, to systematically assess potential impacts on mosquito performance and survival under increasingly natural conditions (Gato et al., 2014, 2021). However, a critical gap remains: reliable estimation of key ecological parameters of released sterile males (such as dispersal and survival) in real-life urban settings, which are essential to the design and scheduling of release strategies.
@@
MRR experiments, combined with spatio-temporal dispersal models, have long been used to infer demographic parameters including survival rates, longevity, and emigration (Lebreton et al., 1992; Silver, 2007; Cordero-Rivera and Stoks, 2008; Grosbois and Gimenez, 2010). Yet, mosquito MRRs pose distinctive challenges: short lifespans and low recapture rates yield sparse, noisy data (Garcia et al., 2016), and classical population-scale models can be inadequate. This motivates the use of microscopic individual-based models (IBMs), which represent each mosquito explicitly, but IBMs are computationally intensive and their stochasticity complicates parameter inference (Hartig et al., 2011).
Our methodological contribution is to integrate two complementary modeling frameworks. At the microscopic level, we model mosquito movement as a two-dimensional Itô diffusion with spatially varying mobility, and we treat death and capture as stochastic events. At the macroscopic level, we derive a reaction–diffusion formulation that is the deterministic limit of the IBM and yields the expected population density and expected captures. This theoretically grounded link allows us to estimate IBM parameters without direct large-scale simulation: inference proceeds by solving the reaction–diffusion system and embedding it in a mechanistic–statistical observation model. Such mechanistic–statistical approaches have proved effective for reaction–diffusion systems under sparse, noisy observations—either when models are constructed from biological principles (Roques et al., 2011; Roques and Bonnefon, 2016) or when they summarize underlying stochastic processes (Roques et al., 2016; Roques et al., 2022)—and naturally fit within a state-space framework (Patterson et al., 2008; Durbin and Koopman, 2012). A central focus here is the explicit modeling of trapping and the joint identifiability of movement, survival, and trap efficiency.
@@
We remark that obtaining good estimates on how mosquito spatial diffusion depends on the heterogeneity of the terrain is important to be able to calibrate mathematical models that take into account spatially heterogeneous environments in order to control mosquito populations in an efficient and robust way (see, for instance, Agbo Bidi et al., 2025).
@@ 2.1 Mark-release-recapture data
The mark-release-recapture (MRR) experiment described in this paper was conducted at a time when wild A. aegypti mosquito populations at the study site exhibited low abundance due to a preceding SIT pilot study (Gato et al., 2021).
@@
Study Design: The study was designed following the guidelines described by Bouyer et al. (2024). Four releases of 10,000 marked mosquitoes each were carried out over consecutive weeks at a single location, referred to as the release point, at the center of the study site. Each week, the mosquitoes were marked with a different color using a fluorescent powder: yellow, red, blue, and pink, respectively. An adult mosquito trap network was deployed to recapture the released mosquitoes and monitor the wild population. The network consisted of twenty-one BG-Sentinel traps baited with BG-lures (Biogents, Germany), distributed along concentric rings at distances of approximately 50, 100, 150, 200, 250, 300 and 400 meters. The traps were installed indoors, at ground level, in the quietest areas of inhabited buildings, and were checked daily for six weeks. The collected biological material was transported to the laboratory (at IPK) in plastic containers, where insects were killed by freezing at -20°C. Mosquitoes were then identified morphologically by species and sex under a stereomicroscope. Males were further classified as either wild, unmarked individuals or marked with different dust colors, using ultraviolet light for identification.
Mosquitoes: A. aegypti were colonized from eggs collected in ovitraps at the study site in 2018. Adults were reared in cages at a density of one mosquito/cm², with a 1:2 male-to-female ratio, and fed a 10% honey solution. The colony was maintained at 28 ± 2°C and 80 ± 10% relative humidity. Porcine blood at 38°C was provided in collagen casings (Fibran, Girona, Spain) once a week for female feeding. Eggs were collected 2–3 days post female feeding and allowed to mature in a wet environment for 3 days. Larvae were reared in 100 cm × 60 cm × 3 cm trays, each containing 4 L of deionized water, at a density of 2 larvae/mL. Larval density was determined using egg quantity-weight regression curves as described in (Zheng et al., 2015). Simultaneous hatching was induced by immersing eggs in 36°C de-oxygenated water under vacuum. First-instar larvae were then transferred to trays. The IAEA standard diet (50% tuna meal, 36% bovine liver powder, and 14% brewer’s yeast) was provided daily at the rates of 0.2, 0.4, 0.8, and 0.6 mg/larva for larval instars I, II, III, and IV, respectively (Puggioli et al., 2013). After the onset of pupation, immatures were collected and sorted by sex using a Fay-Morlan apparatus (John W. Hock Co, Gainesville, FL, USA). Male pupae were dosed by volume into 10 mL plastic tubes, each containing approximately 500 individuals. Batches of 5,000 male pupae were kept in 1 L flat tissue-culture flasks (Thermo Fisher, Waltham, MA, USA) filled with 250 mL of dechlorinated water until the optimal age for irradiation was reached. The flasks were placed horizontally to maximize the water surface area.
@@
Marking: The boxes containing two-day-old adult sterile males were placed in a fridge at 4°C for 15 minutes to immobilize the mosquitoes. The immobilized mosquitoes were transferred in batches of approximately 2,500 specimens to 1-liter plastic containers containing 10.4 mg of fluorescent powder (DayGlo® Color Corp., USA) following previously described procedures (Bouyer et al., 2024). The containers were gently rotated for ten seconds to ensure all mosquitoes came into contact with the pigment. Marked mosquitoes were then transferred to 30 x 30 x 30 cm metallic cages (BioQuip, USA). Mosquitoes were provided with water and honey solution for 24 hours prior to release.
@@ 2.2 Mechanistic models and simulated data
where $dB_{t}$ denotes an increment in time of a standard 2-dimensional Brownian motion (Gardiner, 2009), and $\sigma$ is a Lipschitz-continuous function on $\mathbb{R}^{2}$ that describes the local mosquito mobility, which may vary depending on local conditions.
@@
First, ignoring the dead or alive status of the mosquitoes, the probability density $v(t,x)$ of the random variable $X_{t}$ is given by the standard Fokker-Planck equation (Gardiner, 2009):
@@
Let $p(t,x)$ denote the probability density function of the mosquito being alive at time $t$ and located at $x$. The killing mechanism removes probability mass at a rate $F(x)p(t,x)$. Thus, the corresponding Fokker-Planck equation is given by (see, e.g., Roques et al., 2022):
@@ 2.4 Numerical and software aspects
The 2D MRR simulator is implemented in Python and builds upon the MSE software , which is dedicated to the simulation and calibration of PDE models. The MRR simulator is available in a dedicated Git repository, here [https://forge.inrae.fr/msegrp/mrr-simulator]. Since MSE is distributed through GUIX (Courtès et al., 2024), this work can be reproduced in a controlled environment on Linux operating systems.
@@ 3.2 Parameter estimation using experimental data
To better compare these two models, we use the Akaike Information Criterion (AIC), which introduces a penalty term for the number of parameters in the model (Akaike, 1974): $AIC=-2\ln(\mathcal{L}(\Theta^{*}))+2N_{p},$ where $\mathcal{L}(\Theta^{*})$ is the maximum likelihood of the model, and $N_{p}$ is the number of parameters in the model ($N_{p}=3$ in the homogeneous model, $N_{p}=4$ in the heterogeneous model). The AIC balances the goodness of fit (measured by $2\ln(\mathcal{L}(\Theta^{*}))$) and model complexity (penalized by $2N_{p}$). A lower AIC value indicates a better model.
@@ 4 Discussion
Classical analyses of MRR datasets often summarize dispersal and survival empirically from trap returns (e.g., exponential fits for daily survival, recapture-weighted averages for dispersal, Muir and Kay, 1998). In contrast, our approach links movement, mortality, and capture through explicit mechanisms and computes likelihoods from model-predicted captures at each trap and day, which avoids ad hoc transformations and, crucially, clarifies the interpretation of dispersal metrics. Our estimates are consistent with those of Gato et al. (2021) for lifespan: in that other Cuban field trial, the probability of daily survival (PDS) was obtained by fitting an exponential model to recapture counts, and the average post-release life expectancy was computed as $1/(-\ln\mathrm{PDS})$, yielding 3.76 days; combined with the $\sim$3-day age at release, this corresponds to a total lifespan close to 7 days, in good agreement with our inference (about 8 days). For context, published field MRR studies of wild (non-sterile) A. aegypti males report daily survival probabilities around 0.570.570.700.70 in northern Australia (Muir and Kay, 1998) (implying mean life expectancy of $\approx 1.8$$2.8$ days) and $\approx 0.77$ on the Kenyan coast (McDonald, 1977a) (implying $\approx 3.8$ days). Regarding pre-release age, Muir and Kay (1998) report that released males were 1–2 days post-emergence, whereas McDonald (1977a) describes “newly emerged” males without specifying an exact number of days. These values place our sterile-male post-release estimate ($5$ days) and total lifespan (8 days) slightly above these wild-male estimates.
@@
By contrast, dispersal metrics differ substantially. In (Gato et al., 2021), dispersal was summarized from the locations of recaptured marked males in a BG-Sentinel network arranged in concentric rings up to 400 m, as a mean distance of recaptures (77.3 m) and as flight ranges (43.2 m for 50% of recaptures, 110.5 m for 90%). Methodologically, this “mean distance” is the average of trap-to-release distances weighted by how many mosquitoes were recaptured in each trap; in other words, it is the mean distance conditional on being captured (this is a standard approach, also used in (McDonald, 1977b; Muir and Kay, 1998)). Such a quantity is generally a downward-biased proxy for the population mean displacement because (i) capture probability typically declines with distance as trap density per unit circumference decreases with increasing ring radius, although this effect can be corrected for by applying a weighting factor, as proposed by Muir and Kay (1998); (ii) long-distance movements are right-truncated by the outer trap radius (400 m); and (iii) capture is an absorbing event that truncates trajectories—individuals caught early cannot contribute longer displacements they might otherwise have achieved. For these reasons, the reported distances should be interpreted as lower bounds on true dispersal. In our mechanistic–statistical framework, dispersal is inferred through an explicit diffusion parameter ($\sigma$), independent of trap placement. With the homogeneous-model estimate $\sigma=64~\text{m}/\sqrt{\text{day}}$, the expected distance from the release point is $E(\|X_{t}\|)=\sigma\sqrt{\pi t/2}\approx 80\sqrt{t}$ m, i.e., roughly 180 m after 5 days, naturally exceeding the uncorrected mean of recaptures and aligning with the expectation that trap-weighted averages under-estimate long-range movement.
@@
Regarding potential improvements to the MRR experiment, a promising alternative would be to perform simultaneous releases of mosquitoes marked with distinct colors at different locations, rather than successive weekly releases from the same site. In such a design, frequency data (i.e., the relative proportions of each color captured per trap) could be analyzed instead of absolute abundances, as proposed in (Roques et al., 2016). Using proportions has the advantage of reducing sensitivity to heterogeneity in trapping efficiency across sites (e.g., spatial variations in $\gamma$ or $R$), thereby mitigating potential biases in parameter estimation.